OpenStudy (anonymous):

Let r = 1 + cos θ. Find the horizontal and vertical tangent lines to the polar curve.

OpenStudy (tkhunny):

\(x = r(\theta)\cdot \cos(\theta).\;\;Find\;\dfrac{dx}{d\theta}.\) \(y = r(\theta)\cdot \sin(\theta).\;\;Find\;\dfrac{dy}{d\theta}.\)